### v1

# VisNumBench: Evaluating Number Sense of Multimodal Large Language Models

Abstract
Can Multimodal Large Language Models (MLLMs) develop an intuitive number sense similar to humans? Targeting this problem, we introduce Visual Number Benchmark (VisNumBench) to evaluate the number sense abilities of MLLMs across a wide range of visual numerical tasks. VisNumBench consists of about multiple-choice question-answer pairs derived from both synthetic and real-world visual data, covering seven visual numerical attributes and four types of visual numerical estimation tasks. Our experiments on VisNumBench led to the following key findings: (i) The 17 MLLMs we tested—including open-source models such as Qwen2.5-VL and InternVL2.5, as well as proprietary models like GPT-4o and Gemini 2.0 Flash—perform significantly below human levels in number sense-related tasks. (ii) Multimodal mathematical models and multimodal chain-of-thought (CoT) models did not exhibit significant improvements in number sense abilities. (iii) Stronger MLLMs with larger parameter sizes and broader general abilities demonstrate modest gains in number sense abilities. We believe VisNumBench will serve as a valuable resource for the research community, encouraging further advancements in enhancing MLLMs’ number sense abilities. All benchmark resources, including code and datasets, will be publicly available at https://wwwtttjjj.github.io/VisNumBench/.
1 Introduction
Number sense is an innate cognitive ability shared by both humans and animals through the approximate number system [14]. It enables individuals to perceive, process, and manipulate numerical information intuitively. By fostering a deeper understanding of abstract number concepts, it facilitates the grasp of complex mathematical theories and their practical application in real-world scenarios. Figure 1 illustrates the human ability to perceive and estimate numerical quantities. For instance, a person can quickly recognize a group of five bananas and intuitively estimate that there are about two more groups of the same size. By leveraging this innate number sense and grouping strategy, one can infer that the total number of bananas is approximately fifteen.
Multimodal Large Language Models (MLLMs) have made remarkable strides in tackling complex multimodal tasks [2, 7, 24, 26]. Recent research has focused on enhancing their mathematical and scientific reasoning capabilities by incorporating external tools [31, 49]. To assess these abilities, numerous benchmarks [30, 10, 32, 29, 20, 18] have been developed to evaluate the performance of MLLMs on mathematical reasoning and numerical interpretation tasks. While existing benchmarks effectively assess structured numerical reasoning problems, they primarily emphasize abstract symbolic computation, mathematical problem-solving, or interpreting numerical data in textual contexts. However, these evaluations overlook a critical aspect of human-like numerical cognition: intuitive number sense. Unlike humans, who effortlessly estimate quantities, perceive proportions, and grasp numerical relationships at a glance, MLLMs often depend on explicit reasoning steps rather than perceptual intuition. This limitation raises fundamental questions about whether current models genuinely comprehend numerical concepts or merely manipulate them based on learned patterns in text and images.
VNS-Bench | Angle | Length | Scale | Quantity | Depth | Area | Volume | Total |
VNS-Bench-Synthetic | 170 | 181 | 140 | 196 | 135 | 189 | - | 1011 |
VNS-Bench-Real | 149 | 162 | 143 | 147 | 154 | - | 147 | 902 |
Answer Format | 4/5 options | 3/4 options | 4 options | 3/4 options | 4 options | 4/5 options | 3/4 options | 3/4/5 options |
In this work, we introduce the Visual Number Benchmark (VisNumBench), inspired by human number sense abilities. As illustrated in Figure 2, VisNumBench is structured into two components based on different visual scenarios: VisNumBench-Synthetic and VisNumBench-Real. VisNumBench-Synthetic comprises controlled, synthetic images in which numerical relationships are explicitly defined. VisNumBench-Real contains real-world images, providing a more complex and less controlled environment. Grounded in human number sense, VisNumBench targets seven key dimensions of visual numerical attributes through four distinct types of visual numerical estimation tasks, as depicted in Figure 3.
We evaluated 17 MLLMs on VisNumBench and found that even state-of-the-art models perform poorly on our proposed benchmark, which is shown in Figure 4. Furthermore, our experiments reveal that adopting multimodal mathematical models and multimodal Chain-of-Thought (CoT) models did not lead to substantial performance improvements. However, the performance of the latest models is better than the previous models from the same family. For example, Qwen2.5VL [41] performs better than Qwen2VL [44]. It seems that optimization on data, training techniques, and model architecture will help models improve number sense ability. Within this work, we aim at advancing MLLMs toward higher levels of intelligence by developing models that enhance visual number sense abilities.
The main contributions of this paper are listed as follows:
1.
We introduce VisNumBench, a comprehensive benchmark that integrates diverse data sources and an automated evaluation framework to assess the numerical sense abilities of MLLMs across various visual numeric tasks.
2.
We conduct a comprehensive evaluation of various MLLMs on VisNumBench and find that even the most advanced models still demonstrate limited numerical sense abilities.
3.
Further experiments on historical models from the same family show that their numerical sense abilities have improved over time. To enhance this ability within a short period, more specialized optimizations in data, training techniques, and model architecture may be required.
2 Related Work
2.1 Multimodal Large Language Models
Recent advancements in MLLMs have demonstrated exceptional capabilities across a wide range of tasks. Leveraging multimodal pre-training, MLLMs have achieved outstanding performance in both open-source models [51, 1, 45, 4, 5] and proprietary models [34, 36, 3]. Consequently, these models have been widely adopted in various domains, including mathematical reasoning [55], chart understanding [33], medical image analysis [23], and text-rich image comprehension [56]. Their growing success has spurred the development of an increasing number of benchmarks to assess performance across diverse visual and linguistic tasks.
2.2 Benchmarks for MLLMs
Advancements in MLLMs have led to the development of numerous benchmarks aimed at evaluating model performance across a broad spectrum of general multimodal tasks. Several recent studies [15, 53, 28, 22, 21, 52] have introduced more comprehensive multimodal benchmarks that provide extensive and holistic assessments. Beyond general multimodal tasks, specialized benchmarks have been created to evaluate the mathematical reasoning capabilities of MLLMs. Tasks such as visual reasoning with numbers, arithmetic problem-solving, and algebraic manipulation play a crucial role in assessing both the numerical proficiency and higher-order cognitive abilities of MLLMs. Prominent benchmarks, including MATHVISTA [29], Math-Vision [43], MathOdyssey [13], and SMART-840 [9], are designed to test models on a diverse range of mathematical challenges, such as word problems, equation-solving, and complex multi-step reasoning.
These benchmarks aim to assess the ability of models to understand and process mathematical content in both images and text, as well as their ability to apply mathematical operations in reasoning contexts. However, evaluating MLLMs is crucial not only for measuring their proficiency in traditional mathematical reasoning but also for understanding their ability to handle more tangible, real-world dependent mathematical perception tasks. Humans typically acquire basic mathematical knowledge through intuitive number sense, which they then apply to real-world scenarios. This intuitive understanding of numbers and their relationships is a fundamental aspect of human cognition, enabling the seamless application of mathematical concepts in everyday life. While MLLMs can process complex mathematical problems, they may struggle with tasks that require this intuitive number sense. Therefore, existing benchmarks should not only evaluate the proficiency of models in traditional mathematical reasoning but also assess their ability to apply mathematical concepts in real-world contexts. This would help bridge the gap between abstract problem-solving and practical application.
2.3 Number Sense of MLLMs
In the context of MLLMs, previous research [25, 46, 11] has focused on tasks that rely on ordinal regression to evaluate number sense. Examples of such tasks include Age Estimation [37], Historical Image Dating [35], and Image Aesthetics Assessment [38]. These studies typically focus on estimating or ranking numerical attributes based on visual input, such as predicting the age of a person in an image or determining the historical context of a photograph. While these tasks assess the ability of the model to interpret numerical cues and sequences, they often focus on specific domains and may overlook the broader, more generalizable concept of number sense.
Assessing the number sense of MLLMs refers to evaluating their number sense abilities in a variety of scenarios. This includes tasks such as interpreting measurements, performing approximate calculations, comparing quantities, and identifying numerical relationships in different contexts. Such evaluations provide a better understanding of how models perform on tasks that require flexible and contextual reasoning about numbers while also enhancing their broad applicability and intelligence.
3 VisNumBench
3.1 Overview
We introduce VisNumBench, a benchmark specifically designed to directly evaluate the intuitive numerical abilities of MLLMs. Each instance in VisNumBench comprises a figure, a multiple-choice question, and a corresponding answer label. The dataset statistics of VisNumBench are presented in Table 1.
Compared with previous benchmarks, VisNumBench has the following novel features:
Comprehensive Scenario Integration: VisNumBench incorporates both controlled synthetic figures and intricate real-world scenes, enabling a thorough evaluation of the number sense abilities of MLLMs.
Multidimensional Visual Numerical Attributes: VisNumBench encompasses seven fundamental aspects of number sense—angle, length, scale, quantity, depth, area, and volume—ensuring a rigorous and comprehensive evaluation of the numerical capabilities of MLLMs.
Comprehensive Visual Numerical Estimation Tasks: VisNumBench encompasses four distinct modes of visual numerical estimation—value comparison, value estimation, range estimation, and multiplicative estimation. These diverse tasks enable a thorough evaluation of MLLMs’ ability to estimate numerical values across different visual numerical categories.
Angle | Length | Scale | Quantity | Depth | Area | Average | |
Random | 24.44 | 25.41 | 25.00 | 25.00 | 25.00 | 23.68 | 24.76 |
Open-source MLLMs | |||||||
Phi-3.5-vision | 19.41 | 40.88 | 41.43 | 26.53 | 26.67 | 39.15 | 32.34 |
LLaVA-v1.5-7B | 31.18 | 30.39 | 34.29 | 33.16 | 26.67 | 21.16 | 29.38 |
LLaVA-v1.5-13B | 35.88 | 30.94 | 32.14 | 36.73 | 33.33 | 24.34 | 32.15 |
LLaVA-v1.6-34B | 40.00 | 45.30 | 40.00 | 46.94 | 64.44 | 33.33 | 44.31 |
LLaVA-Onevision-7B | 25.88 | 51.38 | 42.86 | 38.78 | 34.81 | 44.44 | 39.96 |
LLaVA-Onevision-72B | 24.71 | 61.33 | 62.14 | 50.00 | 48.15 | 58.73 | 50.84 |
InternVL2.5-8B | 26.47 | 41.99 | 49.29 | 34.69 | 41.48 | 46.03 | 39.66 |
InternVL2.5-38B | 39.41 | 59.67 | 59.29 | 54.08 | 60.74 | 61.38 | 55.59 |
InternVL2.5-78B | 35.29 | 59.67 | 68.57 | 42.86 | 61.48 | 72.49 | 56.18 |
Janus-Pro-7B | 31.76 | 43.65 | 45.71 | 35.71 | 33.33 | 36.51 | 37.69 |
Qwen2.5-VL-3B | 30.00 | 49.17 | 50.71 | 32.14 | 42.22 | 51.85 | 42.43 |
Qwen2.5-VL-7B | 23.53 | 53.59 | 55.00 | 39.29 | 48.89 | 58.20 | 46.19 |
Qwen2.5-VL-72B | 37.06 | 59.67 | 65.00 | 57.65 | 61.48 | 70.37 | 58.46 |
API-based models | |||||||
GPT-4o | 35.29 | 43.09 | 54.29 | 37.24 | 54.07 | 43.39 | 43.72 |
Gemini 1.5 Flash | 26.47 | 47.41 | 44.40 | 26.02 | 23.70 | 41.27 | 33.33 |
Gemini 2.0 Flash | 31.18 | 57.46 | 81.43 | 55.10 | 51.11 | 70.90 | 57.57 |
Gemini 1.5 Pro | 34.12 | 39.23 | 47.14 | 40.82 | 58.52 | 48.15 | 44.02 |
Human | 90.00 | 96.00 | 100.00 | 96.00 | 98.00 | 92.00 | 95.33 |
Angle | Length | Scale | Quantity | Depth | Volume | Average | |
Random | 25.00 | 25.00 | 25.00 | 27.83 | 25.00 | 25.40 | 25.54 |
Open-source MLLMs | |||||||
Phi-3.5-vision | 30.20 | 37.65 | 27.97 | 48.30 | 48.70 | 29.93 | 37.25 |
LLaVA-v1.5-7B | 22.82 | 32.72 | 25.87 | 36.73 | 25.32 | 27.21 | 28.49 |
LLaVA-v1.5-13B | 28.86 | 43.21 | 29.37 | 46.94 | 49.35 | 41.50 | 40.02 |
LLaVA-v1.6-34B | 28.86 | 54.94 | 23.08 | 68.03 | 63.64 | 63.27 | 50.55 |
LLaVA-Onevision-7B | 18.12 | 44.44 | 20.28 | 64.63 | 44.81 | 50.34 | 40.58 |
LLaVA-Onevision-72B | 17.45 | 57.41 | 44.76 | 74.83 | 48.70 | 61.22 | 50.78 |
InternVL2.5-8B | 28.86 | 34.57 | 15.38 | 64.63 | 49.35 | 47.62 | 40.13 |
InternVL2.5-38B | 30.20 | 51.85 | 26.57 | 83.67 | 61.04 | 58.50 | 52.11 |
InternVL2.5-78B | 36.91 | 58.64 | 48.95 | 79.59 | 52.60 | 62.59 | 56.54 |
Janus-Pro-7B | 22.82 | 32.10 | 35.66 | 48.98 | 35.71 | 30.61 | 34.26 |
Qwen2.5-VL-3B | 30.20 | 44.44 | 35.66 | 51.70 | 43.51 | 49.66 | 42.57 |
Qwen2.5-VL-7B | 24.16 | 38.89 | 32.17 | 59.18 | 48.70 | 42.86 | 41.02 |
Qwen2.5-VL-72B | 34.23 | 50.62 | 43.36 | 80.27 | 52.60 | 59.18 | 53.33 |
API-based models | |||||||
GPT-4o | 27.52 | 30.25 | 37.06 | 60.54 | 35.71 | 47.62 | 39.58 |
Gemini 1.5 Flash | 14.77 | 35.80 | 26.57 | 57.14 | 24.68 | 43.54 | 33.70 |
Gemini 2.0 Flash | 38.93 | 48.77 | 74.14 | 81.63 | 46.10 | 51.70 | 56.54 |
Gemini 1.5 Pro | 30.20 | 45.68 | 27.97 | 68.03 | 64.29 | 55.10 | 48.67 |
Human | 96.00 | 100.00 | 100.00 | 98.00 | 96.00 | 94.00 | 97.33 |
3.2 Data Collection Process
3.2.1 Source of Images
The development of VisNumBench required gathering figures from diverse sources, as depicted in part (a) of Figure 5.
Python Program. We developed a series of Python scripts based on Matplotlib111https://matplotlib.org/ to generate figures by randomly sampling parameters, which are also stored for future use. This approach allows precise control over various numerical properties, ensuring a well-balanced data distribution and minimizing potential biases.
Web Collection. To incorporate more natural and diverse visual data, we gathered figures containing numerical information from publicly available web resources [42, 17, 12]. These figures were carefully curated and filtered to ensure relevance and clarity before designing the corresponding questions.
Images Taken by Us. To better reflect real-world conditions, we manually constructed various number sense scenarios and captured images using a camera. These scenes encompassed real-world measurements, physical counting tasks, and estimation challenges under natural lighting and occlusion conditions.
3.2.2 Construction of QA Pairs and Data Quality Control
As shown in parts (b) and (c) in Figure 5, we involve QA pairs for images from different sources. For the figures generated by the Python program, we manually design different questions based on the specific characteristics of each figure and generate corresponding annotations using the parameters saved We can design the question such as: “In the ABCD sections of the pie chart, which section has the largest/smallest area?” Based on the parameters saved, the answer would be “C / D”. For images from other sources, we manually designed questions and annotated them based on different numerical attributes of the images. In addition, we can generate different numerical estimation tasks based on the answer type. For example, for the question: “Which of the following options is a reasonable estimate of the number of cubes in the figure?”, when the answer is “(50, 75)”, it corresponds to a range estimation task. On the contrary, if the answer is “62”, it belongs to the value estimation task.
We employ a combination of automated and manual methods to design distractors. For numerical answers, we generate alternative options that are easily confusable with the correct answer. Some distractors are constructed based on their inherent properties (e.g., areas A, B, and D in part (c) of Figure 5). These distractors are selected to align with human perceptual biases, ensuring that they appear plausible while remaining distinguishable from the correct answer.
To ensure the high quality of VisNumBench, we meticulously reviewed all collected data and filtered out any ambiguous or unclear entries. More details of the data construction process can be found in Appendix B.
4 Experiments
We evaluate 17 well-known MLLMs from 8 model families, including 13 open-source models: Phi-3.5-vision [1], LLaVA-v1.5 (7B, 13B), and LLaVA-v1.6-34B [27], LLaVA-Onevision (7B, 72B) [23], Qwen2.5-VL (3B, 7B, 72B) [41], InternVL2.5 (8B, 38B, 78B) [5], and Janus-Pro-7B [4]. Additionally, we assess 4 proprietary models: GPT-4o [19], Gemini 1.5 Flash, Gemini 2.0 Flash, and Gemini 1.5 Pro [36].
We randomly selected samples ( QA pairs from each category), with sourced from VisNumBench-Synthetic and from VisNumBench-Real. Human evaluators independently answered each question and provided assessments. Accuracy (%) is reported for all experimental results, and all the results are provided in Tables 2 and 3.
4.1 Evaluation Results and Analysis
From Tables 2 and 3, we observe that the performance of MLLMs is not comparable to that of humans. Among the seven types of questions, quantity-related tasks appear to be the easiest, while angle-related tasks are the most difficult. This is possibly because the amount of data available for quantity-related tasks is significantly greater than that for angle-related tasks. By comparing the evaluations on synthetic and real images, we find that the performance of the same model does not exhibit significant variance. Thus, in terms of numerical reasoning ability, both synthetic and real images present similar challenges for existing MLLMs. Moreover, we observe that the best open-sourced model performs comparably to the best closed-source models. More detailed analyses and discussions are provided in the following subsections.
4.1.1 Performance on VisNumBench-Synthetic
Table 2 presents the results for various MLLMs on the VisNumBench-Synthetic dataset. Among the open-source models, Qwen2.5-VL-72B achieves the best performance, with an average accuracy of . InternVL2.5-38B, InternVL2.5-78B, and LLaVA-v1.6-34B also demonstrate strong performance, each achieving either the best or the second-best accuracy in at least two tasks. LLaVA-v1.6-34B attains the highest accuracy in angle and depth estimation; however, its overall average accuracy is only . LLaVA-Onevision-72B also performs well, achieving the highest accuracy in length estimation at . In general, models with larger parameter sizes tend to exhibit superior performance, aligning with the intuition that larger models can better capture complex numerical relationships and fine-grained visual patterns.
In the API-based models, Gemini 2.0 Flash demonstrates the best performance, achieving an average accuracy of . In contrast, GPT-4o and Gemini 1.5 Pro exhibit comparable performance, albeit with lower average accuracies. Gemini 1.5 Flash yields the weakest performance, with an average accuracy of . Notably, certain open-source models perform on par with or even surpass proprietary models, suggesting that the disparity in numerical reasoning capabilities between open-source and closed-source models is minimal.
4.1.2 Performance on VisNumBench-Real
Accuracy on the VisNumBench-Real dataset shows similar trends. InternVL2.5-78B and Gemini 2.0 Flash stand out with an average accuracy of , achieving near-optimal results across multiple tasks, as shown in Table 3. InternVL2.5-38B attained an exceptionally high accuracy of on the quantity task, while LLaVA-v1.6-34B excelled in the volume task, achieving the highest scores. Qwen2.5-VL-72B demonstrated relatively balanced performance across tasks, yielding a suboptimal average accuracy of .
Surprisingly, Gemini 1.5 Pro achieved the highest accuracy in depth estimation, reaching . However, its overall average accuracy remained unsatisfactory. Other proprietary models, such as GPT-4o and Gemini 1.5 Flash, exhibited relatively weaker performance. In general, the accuracy on VisNumBench-Real is lower than that on VisNumBench-Synthetic, likely due to the increased complexity and variability of real-world images.
4.1.3 Performance on Different Visual Numerical Estimation Tasks
As we analyze performance across different numerical estimation tasks, Figure 6 reveals that in the synthetic scenario, Gemini 2.0 Flash and Qwen2.5-VL-72B achieve the highest performance across all sub-tasks, particularly in range estimation, value estimation, and value comparison, where their scores consistently exceed 60. In contrast, GPT-4o exhibits the lowest performance in all tasks, especially in value comparison and multiplicative estimation. In the real-world scenario, most models achieved their best performance in value comparison tasks, which are also the easiest for humans. Although Gemini 2.0 Flash and InternVL2.5-78B continue to perform well in most tasks, their performance in multiplicative estimation has declined compared to the synthetic scenario. Additionally, GPT-4o continues to perform poorly across all tasks, particularly in multiplicative estimation and value comparison, where it falls significantly behind other models.
Notably, in the multiplicative estimation task, LLaVA-v1.6-34B outperforms all other models by a significant margin. This suggests that certain models may be more specialized for specific types of tasks, and further fine-tuning or adjustments could enhance performance across different tasks.
4.2 Further Analysis
How do math-special models perform on the VisNumBench? To investigate the number sense abilities of math-special models, we introduce two multimodal mathematical models: (1) InternVL2-8B-MPO [47], initialized from InternVL2-8B [8] and fine-tuned on the large-scale multimodal reasoning preference dataset MMPR [47], achieving the score of on MathVista; (2) Math-LLaVA-13B [39], initialized from LLaVA-v1.5-13B and fine-tuned on the MathV360K [39] dataset. As shown in Figure 7, InternVL2-8B-MPO achieved a improvement in synthetic scenarios and a increase in real-world scenarios. Its enhancements are task-specific rather than universally effective across different number sense challenges. In contrast, Math-LLaVA-13B exhibited a polarized performance trend: while it improved by on the synthetic dataset, its accuracy declined by in real-world scenarios. This suggests that although the model benefits from training on synthetic data, it struggles to generalize to the complexity and variability of real-world number sense tasks. Relying solely on synthetic data may be insufficient to enhance number sense capabilities in real-world applications. Additional strategies, such as incorporating more diverse real-world training data or refining model architectures, may be necessary to achieve meaningful improvements.
Average (Synthetic) | Average (Real) | |
Qwen2-VL-2B | 31.85 | 24.94 |
Qwen2.5-VL-3B | 42.24 | 42.57 |
Qwen2-VL-7B | 41.25 | 41.91 |
Qwen2.5-VL-7B | 46.19 | 41.02 |
Qwen2-VL-72B | 54.20 | 46.56 |
Qwen2.5-VL-72B | 58.46 | 53.33 |
InternVL2-8B | 39.56 | 39.58 |
InternVL2.5-8B | 39.66 | 40.13 |
InternVL2-40B | 45.50 | 45.12 |
InternVL2.5-38B | 55.59 | 52.11 |
How do the multimodal reasoning models perform? To examine whether reasoning techniques can enhance the number sense abilities of MLLMs, we evaluate two multimodal reasoning models: Llama-3.2V-11B-cot222https://huggingface.co/Xkev/Llama-3.2V-11B-cot [50] and R1-OneVision-7B333https://github.com/Fancy-MLLM/R1-Onevision. Llama-3.2V-11B-cot is trained using LLaVA-o1-100k [50], achieving a performance improvement on MathVista compared to Llama-3.2-11B-Vision-Instruct. R1-OneVision-7B, trained with a rule-based reinforcement learning technique, attains an accuracy of on Mathverse [54]. Accordingly, we assess these models on our benchmark. The results, presented in Figure 7, indicate that neither Llama-3.2V-11B-cot nor R1-OneVision-7B achieved the expected performance gains. On the contrary, their accuracy dropped significantly—except for a modest improvement by Llama-3.2V-11B-cot in synthetic scenarios—especially in real-world settings. These findings suggest that developing reasoning techniques specifically tailored for number sense abilities may be necessary.
What helps improve the performance? To determine the factors contributing to the improvement of number sense ability in MLLMs, we evaluate historical models from the same family over time, specifically the Qwen-VL family and the InternVL family. The results are presented in Table 4. As observed, the performance of the latest models generally surpasses that of their predecessors. By comparing Qwen2-VL [44] with Qwen2.5-VL [41], as well as InternVL2 with InternVL2.5 [6], we observe improvements in several aspects: (1) data scale and quality, (2) a more powerful encoder, (3) model architecture, and (4) training strategy. These findings suggest that further exploration in these directions is essential for enhancing the number sense abilities of MLLMs.
5 Conclusion
In this work, we introduce VisNumBench, a novel benchmark designed to evaluate MLLMs on core number sense abilities that are inadequately addressed by existing evaluation benchmarks. Our assessment of 17 MLLMs uncovers substantial deficiencies in their capacity to demonstrate human-like number sense. Even the most advanced models still demonstrate limited numerical sense abilities. Further experiments on historical models from the same family show that to enhance this ability within a short period, more specialized optimizations in data, training techniques, and model architecture may be required.
6 Acknowledge
This research was supported by the National Natural Science Foundation of China No.62272315.
Appendix A Appendix Outline
Appendix B Details for VisNumBench Construction
Angle The task may involve recognizing angles in both 2D and 3D contexts, such as the angles between intersecting lines or the angle between the viewpoint and an object. The figures for VisNumBench-Synthetic are either generated using Python programs or sourced from VisOnlyQA [20], whereas the images for VisNumBench-Real are either captured by the authors or collected from Google Images [17].
Length Based on synthetic and real scenes, we designed various types of question-answering questions, such as relative length comparison and multiple estimation of line segments, as well as the estimation of length, height, and proportion of different objects, among others. The figures for VisNumBench-Synthetic are either generated by Python programs or sourced from MathVista [29], whereas the figures for VisNumBench-Real are captured by the authors or collected from Google Images [17].
Scale We provide figures illustrating the coordinates of a point in a coordinate system, the time indicated on a clock, and the temperature displayed on a thermometer. The figures for VisNumBench-Synthetic are generated by Python programs, sourced from MathVista [20], while other figures originate from Google Images [17] and ECharts [12]. In contrast, the statistics for VisNumBench-Real are either captured by the authors or collected from Google Images [17].
Quantity Each figure contains a varying number of objects, such as points or triangles in synthetic scenarios, or hot air balloons or pets in real-world scenarios. The figures for VisNumBench-Synthetic are either generated using Python programs or obtained from MathVista [29], whereas the figures for VisNumBench-Real are sourced from Google Images [17] and the ShanghaiTech dataset [57].
Depth We further refine the “Relative Depth” task in BLINK [16] by incorporating additional choice points and introducing new question-answer formats. MLLMs will be presented with images containing objects at varying depths, requiring them to determine the correct depth order or estimate the relative distances between objects.
The figures for VisNumBench-Synthetic are obtained from the WallpapersCraft website [42] or sourced from MathVista [29] and VSLAM-TartanAir [48]. The figures for VisNumBench-Real are sourced from BLINK [16] and the NYU Depth Dataset V2 [40].
Area VisNumBench-Synthetic contains comparisons and estimations of object area sizes for both identical and different shapes, as well as their multiplicative relationships. The figures in VisNumBench-Synthetic are either generated by Python programs or obtained from VisOnlyQA [20].
Volume Objects are presented from different perspectives and in various sizes, requiring MLLMs to infer relative volume sizes and proportions based on visible dimensions and depth. The figures for VisNumBench-Real are obtained through camera capture or sourced from Google Images [17].
More details of data construction are shown in the Tables 5, 6. Figure 8 shows the dataset statistics of VisNumBench based on various visual numerical estimation tasks. Figure 9 shows how to build a QA pair based on images generated by a Python script.
Python Program | Web Collection | Other Dataset | Total | |
Angle | 138 | 0 | 32 | 170 |
Length | 160 | 0 | 21 | 181 |
Scale | 77 | 50 | 13 | 140 |
Quantity | 185 | 0 | 11 | 196 |
Depth | 0 | 70 | 65 | 135 |
Area | 139 | 0 | 50 | 189 |
Total | 699 | 120 | 192 | 1011 |
Image Taken by Us | Web Collection | Other Dataset | Total | |
Angle | 91 | 58 | 0 | 149 |
Length | 144 | 18 | 0 | 162 |
Scale | 21 | 122 | 0 | 143 |
Quantity | 0 | 113 | 34 | 147 |
Depth | 0 | 0 | 154 | 154 |
Volume | 140 | 7 | 0 | 147 |
Total | 396 | 318 | 188 | 902 |
Appendix C Evalution Details
C.1 Model Access
This section details the model access and model parameters (refer to Table 7). The model responses presented in this paper were collected between January 1 and February 28, 2025.
Phi-3.5-vision | microsoft/Phi-3.5-vision-instruct |
LLaVA-v1.5-7B | liuhaotian/llava-v1.5-7b |
LLaVA-v1.5-13B | liuhaotian/llava-v1.5-13b |
LLaVA-v1.6-34B | liuhaotian/llava-v1.6-34b |
LLaVA-Onevision-7B | llava-hf/llava-onevision-qwen2-72b-si-hf |
LLaVA-Onevision-72B | llava-hf/llava-onevision-qwen2-72b-ov-hf |
InternVL2.5-8B | OpenGVLab/InternVL2_5-8B |
InternVL2.5-38B | OpenGVLab/InternVL2_5-38B |
InternVL2.5-78B | OpenGVLab/InternVL2_5-78B |
Janus-Pro-7B | deepseek-ai/Janus-Pro-7B |
Qwen2.5-VL-3B | Qwen/Qwen2.5-VL-3B-Instruct |
Qwen2.5-VL-7B | Qwen/Qwen2.5-VL-7B-Instruct |
Qwen2.5-VL-72B | Qwen/Qwen2.5-VL-72B-Instruct |
GPT-4o | gpt-4o-2024-08-06 |
Gemini 1.5 Flash | gemini-1.5-flash |
Gemini 2.0 Flash | gemini-2.0-flash |
Gemini 1.5 Pro | gemini-1.5-pro-002 |
C.2 Additional Results
This section provides additional results of experiments in Section 4.2.
Tablse 8 and 9 show the results by multimodal mathematical models and multimodal CoT models, corresponding to Figure 7. Table 10 presents the performance of models of varying sizes from the QwenVL and InternVL families, extending the results reported in Table 4. .
Models | Angle | Length | Scale | Quantity | Depth | Area | Average |
multimodal mathematical models | |||||||
Internvl-8B | 28.24 | 49.72 | 55.00 | 28.57 | 31.11 | 46.03 | 39.56 |
InternVL2-8B-MPO | 22.94 | 48.07 | 63.57 | 33.67 | 28.15 | 48.68 | 40.65 |
LLaVA-v1.5-13B | 35.88 | 30.94 | 32.14 | 36.73 | 33.33 | 24.34 | 32.15 |
Math-LLaVA-13B | 31.76 | 45.30 | 28.57 | 39.80 | 37.78 | 30.16 | 35.81 |
multimodal CoT models | |||||||
Llama-VL-3_2-11B | 29.41 | 41.44 | 58.57 | 47.96 | 42.96 | 44.97 | 43.92 |
Llama-3.2V-11B-cot | 28.82 | 46.41 | 40.71 | 56.12 | 38.52 | 57.14 | 45.50 |
Qwen2.5-VL-7B-Instruct | 23.53 | 53.59 | 55.00 | 39.29 | 48.89 | 58.20 | 46.19 |
R1-Onevision-7B | 33.53 | 37.57 | 56.43 | 34.69 | 23.70 | 47.09 | 38.87 |
Angle | Length | Scale | Quantity | Depth | Area | Average | |
multimodal mathematical models | |||||||
Internvl-8B | 30.87 | 36.42 | 29.37 | 71.43 | 30.52 | 39.46 | 39.58 |
InternVL2-8B-MPO | 30.87 | 35.19 | 29.37 | 72.11 | 34.42 | 38.10 | 39.91 |
LLaVA-v1.5-13B | 28.86 | 43.21 | 29.37 | 46.94 | 49.35 | 41.50 | 40.02 |
Math-LLaVA-13B | 20.13 | 35.80 | 23.78 | 45.58 | 37.66 | 35.37 | 33.15 |
multimodal CoT models | |||||||
Llama-VL-3_2-11B | 38.26 | 40.74 | 30.77 | 69.39 | 38.31 | 42.18 | 43.24 |
Llama-3.2V-11B-cot | 27.52 | 38.89 | 23.08 | 64.63 | 31.82 | 44.22 | 38.36 |
Qwen2.5-VL-7B-Instruct | 24.16 | 38.89 | 32.17 | 59.18 | 48.70 | 42.86 | 41.02 |
R1-Onevision-7B | 28.86 | 39.51 | 44.76 | 51.70 | 32.47 | 32.65 | 38.25 |
Angle | Length | Scale | Quantity | Depth | Area/Volume | Average | |
VisNumBench-Systentic | |||||||
Qwen2-VL-2B | 28.24 | 30.39 | 35.00 | 35.20 | 21.48 | 38.10 | 31.85 |
Qwen2-VL-7B | 27.06 | 45.30 | 55.00 | 44.39 | 34.81 | 46.56 | 42.24 |
Qwen2-VL-72B | 32.94 | 57.46 | 63.57 | 54.59 | 58.52 | 59.79 | 54.20 |
Internvl-8B | 28.24 | 49.72 | 55.00 | 28.57 | 31.11 | 46.03 | 39.56 |
Internvl-40B | 23.53 | 58.56 | 57.14 | 37.24 | 37.78 | 58.20 | 45.50 |
Qwen2.5-VL-3B | 30.00 | 49.17 | 50.71 | 32.14 | 42.22 | 51.85 | 42.43 |
Qwen2.5-VL-7B | 23.53 | 53.59 | 55.00 | 39.29 | 48.89 | 58.20 | 46.19 |
Qwen2.5-VL-72B | 37.06 | 59.67 | 65.00 | 57.65 | 61.48 | 70.37 | 58.46 |
InternVL2.5-8B | 26.47 | 41.99 | 49.29 | 34.69 | 41.48 | 46.03 | 39.66 |
InternVL2.5-38B | 39.41 | 59.67 | 59.29 | 54.08 | 60.74 | 61.38 | 55.59 |
VisNumBench-Real | |||||||
Qwen2-VL-2B | 10.74 | 19.75 | 19.58 | 47.62 | 32.47 | 19.73 | 24.94 |
Qwen2-VL-7B | 19.46 | 38.89 | 30.07 | 67.35 | 41.56 | 54.42 | 41.91 |
Qwen2-VL-72B | 21.48 | 45.06 | 37.06 | 74.83 | 48.70 | 52.38 | 46.56 |
Internvl-8B | 30.87 | 36.42 | 29.37 | 71.43 | 30.52 | 39.46 | 39.58 |
Internvl-40B | 30.87 | 50.00 | 28.67 | 72.79 | 35.71 | 52.38 | 45.12 |
Qwen2.5-VL-3B | 30.20 | 44.44 | 35.66 | 51.70 | 43.51 | 49.66 | 42.57 |
Qwen2.5-VL-7B | 24.16 | 38.89 | 32.17 | 59.18 | 48.70 | 42.86 | 41.02 |
Qwen2.5-VL-72B | 34.23 | 50.62 | 43.36 | 80.27 | 52.60 | 59.18 | 53.33 |
InternVL2.5-8B | 28.86 | 34.57 | 15.38 | 64.63 | 49.35 | 47.62 | 40.13 |
InternVL2.5-38B | 30.20 | 51.85 | 26.57 | 83.67 | 61.04 | 58.50 | 52.11 |
C.3 Error Analysis
We take the results of Gemini 2.0 Flash as an example to conduct an error analysis. Based on these results, we randomly selected 10 erroneous instances for each attribute in each scenario and manually analyzed a total of 120 randomly sampled errors across all tasks. We categorize the errors into two types: (1) Errors resulting from the model’s failure to correctly perceive the image (image perception errors). (2) Errors where the model correctly interprets the image and the question but fails to generate the correct numerical answer (numerical intuition errors). Our findings indicate that 28.3% of the errors stem from image perception errors, particularly in scale-related tasks, where the model struggles to recognize the position of pointers. In 71.7% of cases, the model exhibits numerical intuition errors, which manifest in challenges related to depth estimation, angle relationships, and quantity perception. These findings further confirm that the current model indeed lacks robust numerical intuition.

### v2

@@
 We believe VisNumBench will serve as a valuable resource for the research community, encouraging further advancements in enhancing MLLMs’ number sense abilities.
-All benchmark resources, including code and datasets, will be publicly available at https://wwwtttjjj.github.io/VisNumBench/.
+Code and dataset are available at https://wwwtttjjj.github.io/VisNumBench/.
 1 Introduction
@@
 This limitation raises fundamental questions about whether current models genuinely comprehend numerical concepts or merely manipulate them based on learned patterns in text and images.
-VNS-Bench | Angle | Length | Scale | Quantity | Depth | Area | Volume | Total |
-VNS-Bench-Synthetic | 170 | 181 | 140 | 196 | 135 | 189 | - | 1011 |
-VNS-Bench-Real | 149 | 162 | 143 | 147 | 154 | - | 147 | 902 |
+| VisNumBench | Angle | Length | Scale | Quantity | Depth | Area | Volume | Total |
+| VisNumBench-Synthetic | 170 | 181 | 140 | 196 | 135 | 189 | - | 1011 |
+| VisNumBench-Real | 149 | 162 | 143 | 147 | 154 | - | 147 | 902 |
 | Answer Format | 4/5 options | 3/4 options | 4 options | 3/4 options | 4 options | 4/5 options | 3/4 options | 3/4/5 options |
@@
 VisNumBench-Real contains real-world images, providing a more complex and less controlled environment.
-Grounded in human number sense, VisNumBench targets seven key dimensions of visual numerical attributes through four distinct types of visual numerical estimation tasks, as depicted in Figure 3.
+VisNumBench targets seven key dimensions of visual numerical attributes through four distinct types of visual numerical estimation tasks, as depicted in Figure 3.
 We evaluated 17 MLLMs on VisNumBench and found that even state-of-the-art models perform poorly on our proposed benchmark, which is shown in Figure 4.
@@
 For example, Qwen2.5VL [42] performs better than Qwen2VL [45].
-It seems that optimization on data, training techniques, and model architecture will help models improve number sense ability.
-Within this work, we aim at advancing MLLMs toward higher levels of intelligence by developing models that enhance visual number sense abilities.
+It seems that optimization on data, training techniques, and model architecture will help models improve their number sense ability.
+In this work, we aim to advance MLLMs toward higher levels of intelligence by developing models that enhance visual number sense abilities.
 The main contributions of this paper are listed as follows:
 1.
-We introduce VisNumBench, a comprehensive benchmark that integrates diverse data sources and an automated evaluation framework to assess the numerical sense abilities of MLLMs across various visual numeric tasks.
+We introduce VisNumBench, a comprehensive benchmark integrating diverse data sources and an automated evaluation framework to assess the numerical sense abilities of MLLMs across various visual numerical tasks.
 2.
-We conduct a comprehensive evaluation of various MLLMs on VisNumBench and find that even the most advanced models still demonstrate limited numerical sense abilities.
+We conduct a comprehensive evaluation of various MLLMs on VisNumBench, finding that even the most advanced models still exhibit limited numerical sense.
 3.
@@
 Advancements in MLLMs have led to the development of numerous benchmarks aimed at evaluating model performance across a broad spectrum of general multimodal tasks.
-Several recent studies [15, 53, 28, 22, 21, 52] have introduced more comprehensive multimodal benchmarks that provide extensive and holistic assessments.
+Several recent studies [15, 54, 28, 22, 21, 53, 16, 20] have introduced more comprehensive reasoning and perception multimodal benchmarks that provide extensive and holistic assessments.
 Beyond general multimodal tasks, specialized benchmarks have been created to evaluate the mathematical reasoning capabilities of MLLMs.
@@
 This approach allows precise control over various numerical properties, ensuring a well-balanced data distribution and minimizing potential biases.
-Web Collection.
-To incorporate more natural and diverse visual data, we gathered figures containing numerical information from publicly available web resources [42, 17, 12].
-These figures were carefully curated and filtered to ensure relevance and clarity before designing the corresponding questions.
 Images Taken by Us.
@@
 These scenes encompassed real-world measurements, physical counting tasks, and estimation challenges under natural lighting and occlusion conditions.
-3.2.2 Construction of QA Pairs and Data Quality Control
+3.2.2 QA Pair Construction and Quality Control
 As shown in parts (b) and (c) in Figure 5, we involve QA pairs for images from different sources.
-For the figures generated by the Python program, we manually design different questions based on the specific characteristics of each figure and generate corresponding annotations using the parameters saved We can design the question such as: “In the ABCD sections of the pie chart, which section has the largest/smallest area?” Based on the parameters saved, the answer would be “C / D”.
+For the figures generated by the Python program, we manually design different questions based on the specific characteristics of each figure and generate corresponding annotations using the parameters saved.
+We can design the question such as: “In the ABCD sections of the pie chart, which section has the largest/smallest area?” Based on the parameters saved, the answer would be “C / D”.
 For images from other sources, we manually designed questions and annotated them based on different numerical attributes of the images.
@@
 Some distractors are constructed based on their inherent properties (e.g., areas A, B, and D in part (c) of Figure 5).
-These distractors are selected to align with human perceptual biases, ensuring that they appear plausible while remaining distinguishable from the correct answer.
+These distractors are designed to align with human perceptual biases, appearing plausible yet distinguishable from the correct answer.
 To ensure the high quality of VisNumBench, we meticulously reviewed all collected data and filtered out any ambiguous or unclear entries.
@@
 Additionally, we assess 4 proprietary models: GPT-4o [19], Gemini 1.5 Flash, Gemini 2.0 Flash, and Gemini 1.5 Pro [37].
-We randomly selected samples ( QA pairs from each category), with sourced from VisNumBench-Synthetic and from VisNumBench-Real.
+We randomly selected samples ( QA pairs from each numerical attribute), with sourced from VisNumBench-Synthetic and from VisNumBench-Real.
 Human evaluators independently answered each question and provided assessments.
@@
 Among the seven types of questions, quantity-related tasks appear to be the easiest, while angle-related tasks are the most difficult.
-This is possibly because the amount of data available for quantity-related tasks is significantly greater than that for angle-related tasks.
+This is possibly because the amount of training data available for quantity-related tasks is significantly greater than that for angle-related tasks.
 By comparing the evaluations on synthetic and real images, we find that the performance of the same model does not exhibit significant variance.
@@
 InternVL2.5-38B attained an exceptionally high accuracy of on the quantity task, while LLaVA-v1.6-34B excelled in the volume task, achieving the highest scores.
-Qwen2.5-VL-72B demonstrated relatively balanced performance across tasks, yielding a suboptimal average accuracy of .
+Qwen2.5-VL-72B demonstrated relatively balanced performance, yielding a suboptimal average accuracy of .
 Surprisingly, Gemini 1.5 Pro achieved the highest accuracy in depth estimation, reaching .
@@
 4.1.3 Performance on Different Visual Numerical Estimation Tasks
-As we analyze performance across different numerical estimation tasks, Figure 6 reveals that in the synthetic scenario, Gemini 2.0 Flash and Qwen2.5-VL-72B achieve the highest performance across all sub-tasks, particularly in range estimation, value estimation, and value comparison, where their scores consistently exceed 60.
+As we analyze performance across different numerical estimation tasks, Figure 6 reveals that in the synthetic scenario, Gemini 2.0 Flash and Qwen2.5-VL-72B achieve the highest performance across all visual numerical estimation tasks, particularly in range estimation, value estimation, and value comparison, where their accuracies consistently exceed 60%.
 In contrast, GPT-4o exhibits the lowest performance in all tasks, especially in value comparison and multiplicative estimation.
@@
 Notably, in the multiplicative estimation task, LLaVA-v1.6-34B outperforms all other models by a significant margin.
-This suggests that certain models may be more specialized for specific types of tasks, and further fine-tuning or adjustments could enhance performance across different tasks.
+This suggests that certain models may be more specialized for specific types of tasks, and further fine-tuning could enhance performance across different tasks.
 4.2 Further Analysis
 How do math-special models perform on the VisNumBench?
-To investigate the number sense abilities of math-special models, we introduce two multimodal mathematical models: (1) InternVL2-8B-MPO [47], initialized from InternVL2-8B [8] and fine-tuned on the large-scale multimodal reasoning preference dataset MMPR [47], achieving the score of on MathVista; (2) Math-LLaVA-13B [39], initialized from LLaVA-v1.5-13B and fine-tuned on the MathV360K [39] dataset.
+To investigate the number sense abilities of math-special models, we introduce two multimodal mathematical models: (1) InternVL2-8B-MPO [49], initialized from InternVL2-8B [8] and fine-tuned on the large-scale multimodal reasoning preference dataset MMPR [49], achieving an accuracy of 65.65% on MathVista; (2) Math-LLaVA-13B [40], initialized from LLaVA-v1.5-13B and fine-tuned on the MathV360K [40] dataset.
 As shown in Figure 7, InternVL2-8B-MPO achieved a improvement in synthetic scenarios and a increase in real-world scenarios.
@@
 To examine whether reasoning techniques can enhance the number sense abilities of MLLMs, we evaluate two multimodal reasoning models: Llama-3.2V-11B-cot222https://huggingface.co/Xkev/Llama-3.2V-11B-cot [51] and R1-OneVision-7B333https://github.com/Fancy-MLLM/R1-Onevision.
-Llama-3.2V-11B-cot is trained using LLaVA-o1-100k [50], achieving a performance improvement on MathVista compared to Llama-3.2-11B-Vision-Instruct.
+Llama-3.2V-11B-cot is trained using LLaVA-o1-100k [51], achieving a performance improvement on MathVista compared to Llama-3.2-11B-Vision-Instruct [34].
 R1-OneVision-7B, trained with a rule-based reinforcement learning technique, attains an accuracy of on Mathverse [55].
@@
 6 Acknowledge
-This research was supported by the National Natural Science Foundation of China No.62272315.
+This research was supported by the National Natural Science Foundation of China under Grant No. 62272315.
 Appendix A Appendix Outline
-Appendix B Details for VisNumBench Construction
+Appendix B Details of VisNumBench Construction
 Angle The task may involve recognizing angles in both 2D and 3D contexts, such as the angles between intersecting lines or the angle between the viewpoint and an object.
 The figures for VisNumBench-Synthetic are either generated using Python programs or sourced from VisOnlyQA [20], whereas the images for VisNumBench-Real are either captured by the authors or collected from Google Images [17].
-Length Based on synthetic and real scenes, we designed various types of question-answering questions, such as relative length comparison and multiple estimation of line segments, as well as the estimation of length, height, and proportion of different objects, among others.
-The figures for VisNumBench-Synthetic are either generated by Python programs or sourced from MathVista [29], whereas the figures for VisNumBench-Real are captured by the authors or collected from Google Images [17].
+Length Based on both synthetic and real-world scenes, we designed a variety of question-answering tasks, including relative length comparison, multiple segment estimation, and the estimation of object length, height, and proportion, among others.
+The figures in VisNumBench-Synthetic are either generated using Python programs or sourced from MathVista [30], whereas the figures in VisNumBench-Real are captured by the authors or collected from Google Images [17].
 Scale We provide figures illustrating the coordinates of a point in a coordinate system, the time indicated on a clock, and the temperature displayed on a thermometer.
@@
 Figure 9 shows how to build a QA pair based on images generated by a Python script.
+Python-generated images are , and all others are resized with the longer side capped at 500 pixels.
 | Python Program | Web Collection | Other Dataset | Total | |
@@
 C.1 Model Access
-This section details the model access and model parameters (refer to Table 7). The model responses presented in this paper were collected between January 1 and February 28, 2025.
+This section provides details on model access and parameter settings (refer to Table 7).
+The model responses presented in this paper were collected between January 1 and February 28, 2025.
+We set , while all other parameters were kept at their default values.
 | Phi-3.5-vision | microsoft/Phi-3.5-vision-instruct |
@@
 | Gemini 1.5 Pro | gemini-1.5-pro-002 |
-C.2 Additional Results
+C.2 Detail of Evaluation
+Prompt. Table 8 shows the prompts for evaluation. The one below is the prompt with CoT.
+| Prompt |
+| Question: {QUESTION} |
+| Options: {OPTIONS} |
+| Answer the question based on the most likely options. |
+| Provide only the letter corresponding to your choice as the answer (e.g., ‘(a)’, ‘(b)’, ‘(c)’, ‘(d)’, ‘(e)’). |
+| Question: {QUESTION} |
+| Options: {OPTIONS} |
+| Please think and answer the question based on the most likely options. |
+Post-processing. We use InternVL2.5-38B to extract the selected options from the MLLMs response, and the corresponding extracting prompt is referenced from [16], as shown in Figure 10.
+Human Evaluation.
+Two individuals with backgrounds in computer science, who were not involved in the project, independently participated by responding to the questions through a visualization interface.
+No compensation was provided for their participation.
+C.3 Additional Results
 This section provides additional results of experiments in Section 4.2.
-Tablse 8 and 9 show the results by multimodal mathematical models and multimodal CoT models, corresponding to Figure 7.
-Table 10 presents the performance of models of varying sizes from the QwenVL and InternVL families, extending the results reported in Table 4.
-.
+Table 9 presents the improvements introduced by the CoT prompt.
+It can be observed that, except for Gemini 2.0 Flash, which shows a positive gain on VisNumBench-Synthetic, the accuracy of the other models decreases.
+Tables 10 and 11 report the results of multimodal mathematical models and multimodal CoT models, respectively, corresponding to Figure 7.
+Table 12 presents the performance of models of varying sizes from the QwenVL and InternVL families, extending the results shown in Table 4.
+Table 13 summarizes the overall accuracy of various models on the VisNumBench benchmark, covering both synthetic and real subsets.
+| Models | Angle | Length | Scale | Quantity | Depth | Area/Volume | Average |
+| VisNumBench-Systentic | |||||||
+| InternVL2.5-78B | 27.06 | 51.93 | 68.57 | 46.43 | 51.85 | 74.60 | 53.21 |
+| Qwen2.5-VL-72B | 38.24 | 52.49 | 68.57 | 56.63 | 48.89 | 74.07 | 56.68 |
+| Gemini 2.0 Flash | 34.12 | 55.80 | 85.00 | 58.16 | 55.56 | 74.60 | 60.14 |
+| VisNumBench-Real | |||||||
+| InternVL2.5-78B | 30.87 | 59.26 | 58.74 | 78.23 | 48.70 | 55.10 | 55.10 |
+| Qwen2.5-VL-72B | 32.21 | 52.47 | 52.45 | 70.75 | 46.10 | 53.06 | 51.11 |
+| Gemini 2.0 Flash | 33.56 | 49.38 | 72.03 | 80.95 | 41.56 | 59.86 | 55.88 |
 | Models | Angle | Length | Scale | Quantity | Depth | Area | Average |
@@
 | InternVL2.5-38B | 30.20 | 51.85 | 26.57 | 83.67 | 61.04 | 58.50 | 52.11 |
-C.3 Error Analysis
-We take the results of Gemini 2.0 Flash as an example to conduct an error analysis.
-Based on these results, we randomly selected 10 erroneous instances for each attribute in each scenario and manually analyzed a total of 120 randomly sampled errors across all tasks.
-We categorize the errors into two types: (1) Errors resulting from the model’s failure to correctly perceive the image (image perception errors).
-(2) Errors where the model correctly interprets the image and the question but fails to generate the correct numerical answer (numerical intuition errors).
-Our findings indicate that 28.3% of the errors stem from image perception errors, particularly in scale-related tasks, where the model struggles to recognize the position of pointers.
-In 71.7% of cases, the model exhibits numerical intuition errors, which manifest in challenges related to depth estimation, angle relationships, and quantity perception.
-These findings further confirm that the current model indeed lacks robust numerical intuition.
+| Model | Synthetic (1,011) | Real (902) | Overall |
+| Human | 95.33 | 97.33 | 96.27 |
+| Gemini 2.0 Flash | 57.57 | 56.54 | 57.08 |
+| InternVL2.5-78B | 56.18 | 56.54 | 56.35 |
+| Qwen2.5-VL-72B | 58.46 | 53.33 | 56.04 |
+| InternVL2.5-38B | 55.59 | 52.11 | 53.95 |
+| LLaVA-Onevision-72B | 50.84 | 50.78 | 50.81 |
+| Qwen2-VL-72B | 54.20 | 46.56 | 50.60 |
+| LLaVA-v1.6-34B | 44.31 | 50.55 | 47.25 |
+| Gemini 1.5 Pro | 44.02 | 48.67 | 46.21 |
+| InternVL2-40B | 45.50 | 45.12 | 45.32 |
+| Qwen2.5-VL-7B | 46.19 | 41.02 | 43.75 |
+| Llama-VL-3_2-11B | 43.92 | 43.24 | 43.60 |
+| Qwen2.5-VL-3B | 42.43 | 42.57 | 42.50 |
+| Qwen2-VL-7B | 42.24 | 41.91 | 42.08 |
+| Llama-3.2V-11B-cot | 45.50 | 38.36 | 42.13 |
+| GPT-4o | 43.72 | 39.58 | 41.77 |
+| InternVL2-8B-MPO | 40.65 | 39.91 | 40.30 |
+| LLaVA-Onevision-7B | 39.96 | 40.58 | 40.25 |
+| InternVL2.5-8B | 39.66 | 40.13 | 39.88 |
+| InternVL2-8B | 39.56 | 39.58 | 39.57 |
+| R1-Onevision-7B | 38.87 | 38.25 | 38.58 |
+| LLaVA-v1.5-13B | 32.15 | 40.02 | 35.86 |
+| Janus-Pro-7B | 37.69 | 34.26 | 36.07 |
+| Phi-3.5-vision | 32.34 | 37.25 | 34.66 |
+| Math-LLaVA-13B | 35.81 | 33.15 | 34.56 |
+| Gemini 1.5 Flash | 33.33 | 33.70 | 33.50 |
+| LLaVA-v1.5-7B | 29.38 | 28.49 | 28.96 |
+| Qwen2-VL-2B | 31.85 | 24.94 | 28.59 |
+| Random | 24.76 | 25.54 | 25.13 |
+C.4 Error Analysis
+We use the results of Gemini 2.0 Flash as a case study for error analysis.
+Based on these results, we randomly selected 10 erroneous instances for each attribute in each scenario, manually analyzing a total of 120 randomly sampled errors across all tasks.
+We categorize the errors into two types: (1) Errors arising from the model’s failure to accurately perceive the image (image perception errors); (2) Errors in which the model correctly interprets both the image and the question but fails to produce the correct numerical answer (numerical intuition errors).
+Our analysis reveals that of the errors are due to image perception issues, particularly in scale-related tasks, where the model struggles to identify the positions of pointers.
+The remaining are attributed to numerical intuition errors, which commonly involve difficulties with depth estimation, angle relationships, and quantity perception.
+These findings further substantiate that current models indeed lack robust numerical intuition.
+Appendix D Example Data and Model Outputs